in Test Pause Test Quiz: Chapter 6 Question 4· 1 of 24 Step 1 of 1 ACKSON AUST Suppose SAT Writing scores are normally distributed with a mean of 496 and a standard deviati to students whose number, if necessary. scores are in the top 9 % . What is the minimum score required for the scholarship? Round your answer to the nearest whole 02:14:28 of 111. A university plans to award scholarships Answer2 Points l Keypad Tables
Answer #1

We are trying to find a Zcritical above which the probability in a standard normal distribution is greater than 9%

P(Z > Zcritical) = 0.09

From standard normal distribution table,  Zcritical = 1.35

Thus, X = mean + (std. dev.) * 1.35 = 496 + 1.35*111 = 496 + 149.85 = 645.85

Thus, 645.85 is the minimum marks that will put anyone in the top 9%